OpenStudy (anonymous):

PLEASE HELP! Solve the following system by using graph paper or graphing technology. 3x – 2y = 7 9x – 6y = 3 I need to put these into slope-intercept form so I can solve the equation I am doing, this is the only part I don't know how to do. Thank you!

OpenStudy (anonymous):

solve each equation for y. then just graph it in y=mx+b mode and you can tell where they intersect

OpenStudy (anonymous):

these are parallel lines .parallel lines do not intersects (OR intersects at infinity) slope intercept for is y=(3x-7)/2 y=3/2x-7/2 foe second y=(9x-3)/6 y=9/6x-3/6 y=2/2x-1/2